# Capturing Turbulence with Numerical Dissipation: a Simple Dynamical Model for Unresolved Turbulence in Hydrodynamic Simulations

### v2

@@ 4.2 Distribution of small-scale turbulence
Figure 9: Cell-by-cell distribution of predicted the distribution of total velocity dispersion including subgrid turbulence and thermal motions, $\sigma_{\rm tot}\equiv\sqrt{\sigma_{\rm turb}^{2}+c_{\rm s}^{2}}$, as a function of gas density, colored by the galactocentric radius of the cell. The panels correspond to the simulations with our new semi-implicit LES (top) and explicit LES models (bottom). The diagonal dotted lines show the constant values of the local virial parameter, from top to bottom: $\alpha_{\rm vir}=100$, 10, and 1. The contours show 68%, 95%, and 99% of cells weighted by mass. The data is stacked using 21 snapshots equally spaced between $t=500$ and $700\;{\rm Myr}$. The inner 500 pc, where the effects of shear are strong, are excluded from the plot. In the average ISM, both models produce similar distributions.
@@
Figure 8 shows the face-on view of our simulated galaxy with our new semi-implicit LES (top row) and explicit LES (bottom panel) models for unresolved turbulence. Qualitatively, both models produce similar global ISM structures. The two noticeable differences in the distribution of subgrid turbulent velocity (second column) are the higher level of turbulence on the outskirts and in the center with the SLES model compared to the explicit LES. Our resimulation of the LES run without shear filtering (not shown here) the level of turbulence in the outskirts and the center increases, indicating that this difference is, at least partially, produced by the shear due to differential rotation. Despite these differences, both models produce visually similar distributions of locally variable $\epsilon_{\rm ff}$ (third column) and the resulting distribution of young stars ($<100\;{\rm Myr}$; rightmost column).
@@
The distribution of star formation efficiencies and rates in the two models are similar because the local distribution of total velocity dispersion and gas density—which together define $\alpha_{\rm vir}$ and $\epsilon_{\rm ff}$—are similar. This is explicitly shown in Figure 9. In particular, the dotted lines in the plot indicate constant values of viral parameters, from top to bottom, $\alpha_{\rm vir}=100$, 10, and 1, corresponding to $\epsilon_{\rm ff}\sim 10^{-8}$, 0.01, and 0.23, respectively. Thus, it is particularly interesting that the distributions of $\sigma_{\rm tot}$ are similar in the dense ($n>10\;{\rm cm^{-3}}$) gas because these are the densities at which most of the star formation occurs and where $\sigma_{\rm turb}>c_{\rm s}\sim 1\;{\rm km\;s^{-1}}$.
One of the physical reasons why these distributions are similar in dense regions is that they are largely shaped by compressive heating (via the $PdV$ work) and decay of turbulence (semenov16; semenov17), and these processes that are modeled in exactly the same way in both models (see Section 2). The evolution of gas in this part of the $n$$\sigma_{\rm tot}$ plane is, in fact, analogous to the evolution of gas in our supersonic developed turbulence tests in Figure 5: at high $\sigma_{\rm tot}$, gas density on average increases, implying $\boldsymbol{\nabla}\boldsymbol{\cdot}{\boldsymbol{u}}<0$, while at low $\sigma_{\rm tot}$, gas density decreases and $\boldsymbol{\nabla}\boldsymbol{\cdot}{\boldsymbol{u}}>0$, producing clock-wise circulation of gas in that plane (see Figures 4 and 6 in semenov17). Thus, as long as subgrid turbulence is produced at a reasonable level in diffuse gas, gas compression during the formation of dense regions will bring it to the values set by the interplay between the $PdV$ work and turbulence decay.
In diffuse gas at the disk outskirts, in contrast, the values of $\sigma_{\rm turb}$ are different (recall Figure 8). However, in these regions, turbulence is strongly subsonic, and $\sigma_{\rm tot}$ is dominated by the thermal sound speed, which creates the lower envelope of the $n$$\sigma_{\rm tot}$ distribution at $n<10\;{\rm cm^{-3}}$. As a result, at these densities, small-scale turbulence does not affect gas dynamics, as its contribution to the total pressure is small.
@@
The inner 500 pc was excluded from the plot, as the strong shear in that region produces elevated $\sigma_{\rm turb}$ in the SLES model analogous to the artifact discussed in semenov22. The model behavior in the presence of strong shear as in the galaxy center should be treated with caution (see Section 5), in particular, due to the effects of stellar feedback associated with the differences in the SFR in the central region associated with such elevated $\sigma_{\rm turb}$. In addition, the gas behavior in the center of $L_{\star}$ galaxies is expected to be strongly affected by active galactic nuclei, not modeled in this study. A detailed investigation of these effects requires a separate study.
@@ 4.3 Global and kiloparsec-scale star formation
Figure 10: Evolution of global star formation properties in our galaxy simulations with two different models for unresolved turbulence. The panels show, from top to bottom: the total SFR, the total mass of star-forming gas, its average density, and local star formation efficiency per freefall time, $\epsilon_{\rm ff}$. Solid lines show the tracks smoothed on a 100 Myr timescale using a Gaussian filter. The SLES model results in $\sim 2$ times lower SFR than the explicit LES simulation. This difference is mainly due to the difference in the predicted local $\epsilon_{\rm ff}$, while the total mass and average density of star-forming gas are similar between the runs.
@@
Figure 10 shows the evolution of galaxy properties describing global star formation in our simulations: the global SFR ($\dot{M}_{\rm\star}$), the total mass of star-forming gas ($M_{\rm sf}$), its average density ($\overline{n}_{\rm sf}$), and SFR-weighted average local star-formation efficiency per freefall time ($\langle\epsilon_{\rm ff}\rangle$). The differences in the global SFR can be explained via the differences in the other three quantities as the SFR can be expressed as
@@
where $\overline{t}_{\rm ff}$ is the inverse of $\langle t_{\rm ff}^{-1}\rangle$ weighted by $\epsilon_{\rm ff}\,\rho\,dV$ for consistency with Equation (9), and $\overline{n}_{\rm sf}$ is the density corresponding to this freefall time. [^5] This is one of the weighting schemes that makes Equation (12) consistent with (9); for an alternative scheme see, e.g., semenov24b.
@@
Figure 11 compares the model predictions on specially resolved kiloparsec scales. Specifically, the figure compares the scaling of the average subgrid velocity dispersion (top) and $\epsilon_{\rm ff}$ (bottom) in star-forming gas as a function of the total gas density in $1\times 1$ kpc patches. The quantities on $y$-axes are averaged over the star-forming gas with $\epsilon_{\rm ff}>0.1\%$ with $\sigma$ being mass-weighted and $\epsilon_{\rm ff}$ being SFR-weighted. In agreement with the above results, the SLES model predicts higher $\sigma$ and lower $\epsilon_{\rm ff}$ in the star-forming gas than the explicit LES model. However, the distributions predicted in both models overlap significantly and are in the ballpark of the values observed in the PHANGS sample of galaxies from (utomo17) and sun20. The low $\epsilon_{\rm ff}\sim 1\%$ values are a generic prediction of this $\alpha_{\rm vir}$-dependent $\epsilon_{\rm ff}$ model in both idealized isolated disk (semenov16; semenov21-tf; polzin24) and cosmological simulations (kretschmer20; segovia-otero24; semenov24b), which results from the regulation of $\alpha_{\rm vir}$ in star-forming gas to a narrow range due to stellar feedback (polzin24). Note that the observed $\epsilon_{\rm ff}$ range is provided in the figure only for reference, and the direct one-to-one comparison with simulation results requires a more careful analysis due to the evolutionary biases induced in the observational estimates of $\epsilon_{\rm ff}$ (e.g., segovia-otero24).
The trend of $\sigma$ and $\epsilon_{\rm ff}$ with $\Sigma_{\rm gas}$ is also interesting by itself. Both $\sigma$ and $\epsilon_{\rm ff}$ increase with $\Sigma_{\rm gas}$. This result is non-trivial because, if $\sigma$ were to increase at fixed local gas density, the virial parameter, $\alpha_{\rm vir}\propto\sigma^{2}/\rho$, would also increase which would cause the decrease of $\epsilon_{\rm ff}$ (see Equations (10) and (11)). Thus, the increasing trend of $\epsilon_{\rm ff}$ with $\Sigma_{\rm gas}$ can be explained by the local gas density of star-forming gas also increasing with $\Sigma_{\rm gas}$ leading to the decrease of the virial parameter.
@@ 5 Discussion
The new Semi-implicit Large Eddy Simulation (SLES) model that we introduced combines features of both explicit and implicit LES. As an explicit LES, it carries the total turbulent energy on unresolved scales as an independent energy variable, enabling one to predict the local turbulent state of gas on unresolved scales. Unlike explicit LES, the source terms describing the turbulent cascade of energy from the resolved to subgrid scales are not modeled explicitly. Instead, the local numerical dissipation, calculated implicitly, is used as a source term.
@@
This approach is motivated by the generic property of a turbulent flow that the dissipation rate is set by the rate of the kinetic energy cascade from the large scales. In simulations where the physical dissipation scale is not resolved, the flow on the scales close to resolution is shaped by the numerical diffusivity of the adopted hydrodynamic method to ensure dissipation of the energy incoming from larger scales, making this diffusivity an effective model for the integral effect of the turbulent motions on unresolved scales. This idea constitutes the basis of the implicit LES (ILES).
@@
Although, at first glance, the fact that numerical errors can conspire to produce a physical result might appear surprising, it has both physical and numerical foundations, which are extensively discussed in boris92; margolin02; grinstein. In short, from the physical standpoint, the turbulent energy spectrum is sufficiently steep so that the small-scale turbulence has only a weak dynamic effect on the resolved flow, alleviating resolution requirements for modeling the largest scales of turbulent flows. At the same time, the spectrum is not too steep, so that the small-scale motions can efficiently mix inhomogeneities produced on larger scales. Except for strong shocks (that we discuss below), the kinetic energy also cascades through all inertial-range scales, making the model applicable as long as the onset of the turbulent cascade is resolved. To satisfy these properties, the numerical method is required to respect the generic properties of the underlying hydrodynamic equations—such as conservation, causality, and lack of spurious oscillations—to ensure that the kinetic energy cascades at the correct rate from large to small scales, where numerical diffusivity dissipates this energy at the same rate.
@@
It is worth noting that both explicit and implicit LES rely on the idea of a local kinetic energy cascade continuous in spatial scales. The notion of cascade becomes nuanced for supersonic turbulence due to compressibility effects. In general, cascades in both kinetic and thermal energies need to be considered in this case as they can exchange energy on any scale as a result of the $PdV$ work done by the gas during compression and expansion (pressure dilatation; see, e.g., aluie13; eyink18 and also garnier for the review of LES models for compressible turbulence). One special case of compressible turbulence for which the presence of a kinetic energy cascade has been derived analytically is the isothermal case (aluie12). For this reason, the above issue is also alleviated in ISM and galaxy formation applications, as most of the volume is typically occupied by the subsonic or transonic, weakly compressible hot ($T>10^{5}$ K) and warm ($T\sim 10^{4}$ K) media, while most of the mass and star formation resides in the cold ($T\lesssim 100$ K) medium which is, albeit supersonic, is nearly isothermal due to strong radiative losses.
@@
The key difference between our semi-implicit method and ILES is that the latter does not explicitly distinguish between subgrid turbulent and thermal energies. Owing to energy conservation, in ILES, all dissipated energy instantly appears as heat, bypassing the unresolved turbulent cascade. In contrast, in SLES we track subgrid turbulent and thermal energies separately and explicitly model the decay of turbulence into heat. For adiabatic flows, SLES and ILES produce identical results as both $e_{\rm turb}$ and $e_{\rm th}$ are described with an ideal gas equation of state with $\gamma=5/3$, making the total pressure independent of the partitioning between $e_{\rm turb}$ and $e_{\rm th}$. In the presence of cooling and heating, however, the results are different because only $e_{\rm th}$ is subject to these processes, while $e_{\rm turb}$ dissipates on a different timescale, close to cell-crossing (see Section 2.4). For example, in dense ISM, $e_{\rm th}$ is subject to strong radiative losses, while $e_{\rm turb}$ can retain some of the dissipated energy, delaying cooling and providing local non-thermal support in the cold ISM. Note that, unlike turbulent viscosity, this turbulent pressure must be explicitly included in hydrodynamic equations analogously with thermal pressure.
@@
The small-scale (close to resolution) structure of the flow and the numerical dissipation scale where the turbulent power spectrum truncates strongly depends on the resolution and the details of the numerical solver (e.g., kitsionas09; kritsuk11). This implies that the effective spatial scale at which the subgrid turbulent energy is estimated in our method—set by the numerical dissipation scale—will vary with the details of the hydrodynamic method, implying that the turbulence decay rate needs to be recalibrated when used in different codes (see Section 3.2). In the context of ILES and SLES, the second-order finite volume methods are particularly attractive because the leading (third) order of their truncation errors matches that of the explicit subgrid terms in LES models (margolin02). The ART code used in this paper is one example of such a second-order finite-volume hydrodynamic code.
@@
In such second-order accurate methods, numerical dissipation starts acting on relatively large scales, $\sim$30 cells and dominates on $\sim$5–10 cells (e.g., kitsionas09; federrath10; federrath11; kritsuk11). As a result, the flow and the distribution of $e_{\rm turb}$ become smoothed by the numerical diffusivity on a similar scale (recall Figure 2). However, the resulting distribution of $e_{\rm turb}$ can still be used to predict the distributions of turbulence properties, such as $\alpha_{\rm vir}$ or Mach numbers. This is because the subgrid model captures the correlations between $e_{\rm turb}$ and other hydrodynamic properties (recall Figure 5), even though they all are affected by numerical diffusivity. The predicted distributions of $\alpha_{\rm vir}$ or Mach numbers, in turn, can be used to improve the subgrid modeling of processes dependent on the turbulent structure of gas. In the context of galaxy formation, these processes include star formation (see Section 4; also, braun15; semenov16; semenov21-tf; kretschmer20), turbulent diffusion of metals (e.g., rennehan21), cooling, heating, and chemical reaction rates, acceleration and transport of cosmic rays, microscopic dynamo (e.g., liu22), among others.
@@
The existing studies of the effects of unresolved turbulence listed above rely on explicit LES models. The key advantage of SLES over explicit models is the simplicity of implementation, as it reuses the functionality present in many hydrodynamic codes used for galaxy formation, specifically, the dual-energy formalism (see Section 2.1; ryu93; bryan95). In this formalism, thermal energy is modeled twice: as a part of conserved total energy and separately, following an adiabatic equation, enabling one to compute gas temperature in strongly supersonic regions where the difference $e_{\rm tot}-e_{\rm kin}=e_{\rm th}\ll e_{\rm tot}$ becomes numerically unreliable. SLES can be implemented by introducing an additional energy variable describing unresolved turbulence and setting it to $e_{\rm turb}=e_{\rm tot}-e_{\rm kin}-e_{\rm th}$ after each step. As $e_{\rm tot}$ is conserved and $e_{\rm th}$ is modeled adiabatically, this step implicitly puts all dissipated energy into $e_{\rm turb}$ (see also semenov22). In principle, one may not evolve $e_{\rm turb}$ explicitly at all and instead apply all appropriate terms to $e_{\rm th}$ and use the above equation to calculate $e_{\rm turb}$ when needed. However, we choose to follow $e_{\rm turb}$ as a separate variable as it enables us to compute it in the regions where the above difference is dominated by numerical errors, by analogy with the original dual-energy formalism.
The SLES model can also be straightforwardly generalized in the presence of magnetic fields. In this case, $e_{\rm turb}$ should be interpreted as total subgrid kinetic plus magnetic energy, while the implicit sourcing by numerical dissipation would encompass the integral effect of LES-like viscosity and resistivity terms entering the MHD equations (see, e.g., grete16; vlaykov16, for an example of such explicit subgrid MHD models).
@@
The results of Section 4 show that the distribution of $e_{\rm turb}$ in cold, supersonic ISM is only weakly sensitive to the choice of the model. This is because in this regime the production of small-scale turbulence is dominated by the heating of pre-existing turbulence via the $PdV$ work exerted by gas undergoing compression in spiral arms, collision of SNe-driven shells, and over disk instabilities (see semenov16; semenov17). Thus, the choice of the model for the cascade term, $\Sigma$, in Equation (1) might not be as crucial for predicting the turbulent properties of the cold ISM, as long as the adiabatic $PdV$ term and turbulence decay, $\epsilon$, are modeled appropriately. However, significant differences do arise in the regions where the cascade of energy from resolved scales dominates the sourcing of $e_{\rm turb}$.
@@
One example is the disk outskirts and very center, where the SLES model predicts higher turbulent velocities than the explicit LES model (Section 4). One reason for this discrepancy is the difference in the treatment of shearing flows. Our explicit model adopts a shear-improved LES closure where the bulk flow is filtered out and only the fluctuating part of the velocity field is interpreted as a part of the turbulent cascade (leveque07, see also Section 2.2). To avoid spurious production of $e_{\rm turb}$ in stable shearing flows, such as differentially rotating disks, one needs to adopt a similar treatment in the SLES model. For example, instead of putting all dissipated energy into $e_{\rm turb}$ one can make it a function of the difference between the local filtered and total flows, such that in purely shearing flows kinetic energy is dissipated directly into heat. We leave the investigation of such models to future work.

### v3

@@ 4.2 Distribution of Small-scale Turbulence
Figure 9: Cell-by-cell distribution of the predicted total velocity dispersion, including subgrid turbulence and thermal motions, $\sigma_{\rm tot}\equiv\sqrt{\sigma_{\rm turb}^{2}+c_{\rm s}^{2}}$, as a function of gas density, colored by the galactocentric radius of the cell. The panels correspond to the simulations with our new SLES (top) and explicit LES models (bottom). The diagonal dotted lines show the constant values of the local virial parameter: from top to bottom, $\alpha_{\rm vir}=100$, 10, and 1. The contours show 68%, 95%, and 99% of cells weighted by mass. The data is stacked using 21 snapshots equally spaced between $t=500$ and $700\;{\rm Myr}$. The inner 500 pc, where the effects of shear are strong, is excluded from the plot. In the average ISM, both models produce similar distributions.
@@
Figure 8 shows the face-on view of our simulated galaxy with our new SLES (top row) and explicit LES (bottom panel) models for unresolved turbulence. Qualitatively, both models produce similar global ISM structures. The two noticeable differences in the distribution of subgrid turbulent velocity (second column) are the higher level of turbulence on the outskirts and in the center with the SLES model compared to the explicit LES. In our resimulation of the LES run without shear filtering (not shown here), the level of turbulence in the outskirts and the center increases, indicating that this difference is, at least partially, produced by the shear due to differential rotation. Despite these differences, both models produce visually similar distributions of locally variable $\epsilon_{\rm ff}$ (third column) and the resulting distribution of young stars ($100\;{\rm Myr}$; rightmost column).
@@
The distributions of star formation efficiencies and rates in the two models are similar because the local distributions of total velocity dispersion and gas density—which together define $\alpha_{\rm vir}$ and $\epsilon_{\rm ff}$—are similar. This is explicitly shown in Figure 9. In particular, the dotted lines in the plot indicate constant values of virial parameters, from top to bottom, $\alpha_{\rm vir}=100$, 10, and 1, corresponding to $\epsilon_{\rm ff}\sim 10^{-8}$, 0.01, and 0.23, respectively. Thus, it is particularly interesting that the distributions of $\sigma_{\rm tot}$ are similar in the dense ($n>10\;{\rm cm^{-3}}$) gas because these are the densities at which most of the star formation occurs and where $\sigma_{\rm turb}>c_{\rm s}\sim 1\;{\rm km\;s^{-1}}$.
[…]
@@
The inner 500 pc is excluded from the plot, as the strong shear in that region produces elevated $\sigma_{\rm turb}$ in the SLES model analogous to the artifact discussed in semenov22. The model behavior in the presence of strong shear, as in the galaxy center, should be treated with caution (see Section 5), in particular, due to the effects of stellar feedback associated with the differences in the SFR in the central region associated with such elevated $\sigma_{\rm turb}$. In addition, the gas behavior in the center of $L_{\star}$ galaxies is expected to be strongly affected by active galactic nuclei, not modeled in this study. A detailed investigation of these effects requires a separate study.
@@ 4.3 Global and Kiloparsec-scale Star Formation
Figure 10: Evolution of global star formation properties in our galaxy simulations with two different models for unresolved turbulence. The panels show, from top to bottom, the total SFR, the total mass of star-forming gas, its average density, and the local star formation efficiency per freefall time, $\epsilon_{\rm ff}$. Solid lines show the tracks smoothed on a 100 Myr timescale using a Gaussian filter. The SLES model results in $\sim 2$ times lower SFR than the explicit LES. This difference is mainly due to the difference in the predicted local $\epsilon_{\rm ff}$, while the total mass and average density of star-forming gas are similar between the runs.
@@
Figure 10 shows the evolution of galaxy properties describing global star formation in our simulations: the global SFR ($\dot{M}_{\rm\star}$), the total mass of star-forming gas ($M_{\rm sf}$), its average density ($\overline{n}_{\rm sf}$), and the SFR-weighted average local star-formation efficiency per freefall time ($\langle\epsilon_{\rm ff}\rangle$). The differences in the global SFR can be explained via the differences in the other three quantities, as the SFR can be expressed as
@@
where $\overline{t}_{\rm ff}$ is the inverse of $\langle t_{\rm ff}^{-1}\rangle$ weighted by $\epsilon_{\rm ff}\,\rho\,dV$ for consistency with Equation (9), and $\overline{n}_{\rm sf}$ is the density corresponding to this freefall time. [^5] This is one of the weighting schemes that make Equation (12) consistent with (9); for an alternative scheme, see, e.g., semenov24b.
@@
Figure 11 compares the model predictions on spatially resolved kiloparsec scales. Specifically, the figure compares the scaling of the average subgrid velocity dispersion (top) and $\epsilon_{\rm ff}$ (bottom) in star-forming gas as a function of the total gas density in $1\times 1$ kpc patches. The quantities on $y$-axes are averaged over the star-forming gas with $\epsilon_{\rm ff}>0.1\%$ with $\sigma$ being mass-weighted and $\epsilon_{\rm ff}$ being SFR-weighted. In agreement with the above results, the SLES model predicts higher $\sigma$ and lower $\epsilon_{\rm ff}$ in the star-forming gas than the explicit LES model. However, the distributions predicted in both models overlap significantly and are in the ballpark of the values observed in the PHANGS sample of galaxies from (utomo17) and sun20. The low $\epsilon_{\rm ff}\sim 1\%$ values are a generic prediction of this $\alpha_{\rm vir}$-dependent $\epsilon_{\rm ff}$ model in both idealized isolated disk (semenov16; semenov21-tf; polzin24) and cosmological simulations (kretschmer20; segovia-otero24; semenov24b), which results from the regulation of $\alpha_{\rm vir}$ in star-forming gas to a narrow range due to stellar feedback (polzin24). Note that the observed $\epsilon_{\rm ff}$ range is provided in the figure only for reference, and the direct one-to-one comparison with simulation results requires a more careful analysis due to the evolutionary biases induced in the observational estimates of $\epsilon_{\rm ff}$ (e.g., segovia-otero24).
The trend of $\sigma$ and $\epsilon_{\rm ff}$ with $\Sigma_{\rm gas}$ is also interesting by itself. Both $\sigma$ and $\epsilon_{\rm ff}$ increase with $\Sigma_{\rm gas}$. This result is nontrivial because, if $\sigma$ were to increase at fixed local gas density, the virial parameter, $\alpha_{\rm vir}\propto\sigma^{2}/\rho$, would also increase, which would cause the decrease in $\epsilon_{\rm ff}$ (see Equations (10) and (11)). Thus, the increasing trend of $\epsilon_{\rm ff}$ with $\Sigma_{\rm gas}$ can be explained by the local gas density of star-forming gas also increasing with $\Sigma_{\rm gas}$, leading to the decrease of the virial parameter.
@@ 5 Discussion
The new SLES model that we introduced combines features of both explicit and implicit LES. As an explicit LES, it carries the total turbulent energy on unresolved scales as an independent energy variable, enabling one to predict the local turbulent state of gas on unresolved scales. Unlike explicit LES, the source terms describing the turbulent cascade of energy from the resolved to subgrid scales are not modeled explicitly. Instead, the local numerical dissipation, calculated implicitly, is used as a source term.
@@
This approach is motivated by the generic property of a turbulent flow that the dissipation rate is set by the rate of the kinetic energy cascade from large scales. In simulations where the physical dissipation scale is not resolved, the flow on the scales close to resolution is shaped by the numerical diffusivity of the adopted hydrodynamic method to ensure dissipation of the energy incoming from larger scales, making this diffusivity an effective model for the integral effect of the turbulent motions on unresolved scales. This idea constitutes the basis of the ILES.
@@
Although at first glance the fact that numerical errors can conspire to produce a physical result might appear surprising, it has both physical and numerical foundations, which are extensively discussed in boris92; margolin02; grinstein. In short, from the physical standpoint, the turbulent energy spectrum is sufficiently steep that the small-scale turbulence has only a weak dynamic effect on the resolved flow, alleviating resolution requirements for modeling the largest scales of turbulent flows. At the same time, the spectrum is not too steep, so that the small-scale motions can efficiently mix inhomogeneities produced on larger scales. Except for strong shocks (which we discuss below), the kinetic energy also cascades through all inertial-range scales, making the model applicable as long as the onset of the turbulent cascade is resolved. To satisfy these properties, the numerical method is required to respect the generic properties of the underlying hydrodynamic equations—such as conservation, causality, and lack of spurious oscillations—to ensure that the kinetic energy cascades at the correct rate from large to small scales, where numerical diffusivity dissipates this energy at the same rate.
@@
It is worth noting that both explicit LES and ILES rely on the idea of a local kinetic energy cascade continuous in spatial scales. The notion of cascade becomes nuanced for supersonic turbulence due to compressibility effects. In general, cascades in both kinetic and thermal energies need to be considered in this case as they can exchange energy on any scale as a result of the $PdV$ work done by the gas during compression and expansion (pressure dilatation; see, e.g., aluie13; eyink18; see also garnier for the review of LES models for compressible turbulence). One special case of compressible turbulence for which the presence of a kinetic energy cascade has been derived analytically is the isothermal case (aluie12). For this reason, the above issue is also alleviated in ISM and galaxy formation applications, as most of the volume is typically occupied by the subsonic or transonic, weakly compressible hot ($T>10^{5}$ K) and warm ($T\sim 10^{4}$ K) media, while most of the mass and star formation resides in the cold ($T\lesssim 100$ K) medium, which, albeit supersonic, is nearly isothermal due to strong radiative losses.
@@
The key difference between our semi-implicit method and ILES is that the latter does not explicitly distinguish between subgrid turbulent and thermal energies. Owing to energy conservation, in ILES, all dissipated energy instantly appears as heat, bypassing the unresolved turbulent cascade. In contrast, in SLES, we track subgrid turbulent and thermal energies separately and explicitly model the decay of turbulence into heat. For adiabatic flows, SLES and ILES produce identical results as both $e_{\rm turb}$ and $e_{\rm th}$ are described with an ideal gas equation of state with $\gamma=5/3$, making the total pressure independent of the partitioning between $e_{\rm turb}$ and $e_{\rm th}$. In the presence of cooling and heating, however, the results are different because only $e_{\rm th}$ is subject to these processes, while $e_{\rm turb}$ dissipates on a different timescale, close to cell-crossing (see Section 2.4). For example, in dense ISM, $e_{\rm th}$ is subject to strong radiative losses, while $e_{\rm turb}$ can retain some of the dissipated energy, delaying cooling and providing local nonthermal support in the cold ISM. Note that, unlike turbulent viscosity, this turbulent pressure must be explicitly included in hydrodynamic equations, analogously with thermal pressure.
@@
The small-scale (close-to-resolution) structure of the flow and the numerical dissipation scale where the turbulent power spectrum truncates strongly depend on the resolution and the details of the numerical solver (e.g., kitsionas09; kritsuk11). This implies that the effective spatial scale at which the subgrid turbulent energy is estimated in our method—set by the numerical dissipation scale—will vary with the details of the hydrodynamic method, implying that the turbulence decay rate needs to be recalibrated when used in different codes (see Section 3.2). In the context of ILES and SLES, the second-order finite-volume methods are particularly attractive because the leading (third) order of their truncation errors matches that of the explicit subgrid terms in LES models (margolin02). The ART code used in this paper is one example of such a second-order finite-volume hydrodynamic code.
In such second-order accurate methods, numerical dissipation starts acting on relatively large scales, $\sim$30 cells, and dominates on $\sim$5–10 cells (e.g., kitsionas09; federrath10; federrath11; kritsuk11). For example, grete23 quantified the resolution dependence of the numerical dissipation scale using energy transfer analysis and reported that dissipation peaks at $\sim$5–10 cells, with a tail extending to scales up to $\sim$5 times larger (see their Figure 4 and Table 1). In the context of the current study, these results imply that the subgrid turbulent energy predicted by the SLES model should be interpreted as the kinetic energy on scales smaller than the dissipation scale, convolved with the code-specific scale dependence of dissipation.
@@
The flow and the distribution of $e_{\rm turb}$ become smoothed by the numerical diffusivity on a similar scale (recall Figure 2). However, the resulting distribution of $e_{\rm turb}$ can still be used to predict the distributions of turbulence properties, such as $\alpha_{\rm vir}$ or Mach numbers. This is because the subgrid model captures the correlations between $e_{\rm turb}$ and other hydrodynamic properties (recall Figure 5), even though they are all affected by numerical diffusivity. The predicted distributions of $\alpha_{\rm vir}$ or Mach numbers, in turn, can be used to improve the subgrid modeling of processes dependent on the turbulent structure of gas. In the context of galaxy formation, these processes include star formation (see Section 4; see also braun15; semenov16; semenov21-tf; kretschmer20); turbulent diffusion of metals (e.g., rennehan21); cooling, heating, and chemical reaction rates; acceleration and transport of cosmic rays; and microscopic dynamo (e.g., liu22), among others.
@@
The existing studies of the effects of unresolved turbulence listed above rely on explicit LES models. The key advantage of SLES over explicit models is the simplicity of implementation, as it reuses the functionality present in many hydrodynamic codes used for galaxy formation, specifically the dual-energy formalism (see Section 2.1; ryu93; bryan95). In this formalism, thermal energy is modeled twice: as a part of conserved total energy and separately, following an adiabatic equation, enabling one to compute gas temperature in strongly supersonic regions where the difference $e_{\rm tot}-e_{\rm kin}=e_{\rm th}\ll e_{\rm tot}$ becomes numerically unreliable. SLES can be implemented by introducing an additional energy variable describing unresolved turbulence and setting it to $e_{\rm turb}=e_{\rm tot}-e_{\rm kin}-e_{\rm th}$ after each step. As $e_{\rm tot}$ is conserved and $e_{\rm th}$ is modeled adiabatically, this step implicitly puts all dissipated energy into $e_{\rm turb}$ (see also semenov22).
As a result, the SLES model imposes negligible costs in terms of both computation and memory. The computational overhead is minimal because the energy cascade term reuses calculations already performed in the code without an LES model (Section 2.3), while all other source and sink terms are either described by analytical functions (e.g., turbulence decay and SN driving) or treated in exactly the same way as in the explicit LES model (e.g., turbulent diffusion; Section 2.4). The memory requirements are also low, since only one additional state variable is needed to track $e_{\rm turb}$ in each cell. In principle, one may not introduce a new variable for $e_{\rm turb}$ at all and instead apply all appropriate terms to $e_{\rm th}$ and use the above equation to calculate $e_{\rm turb}$ when needed. However, we choose to follow $e_{\rm turb}$ as a separate variable as it enables us to compute it in regions where the above difference is dominated by numerical errors, by analogy with the original dual-energy formalism.
The SLES model can potentially be generalized in the presence of magnetic fields. In this case, $e_{\rm turb}$ should be interpreted as the subgrid kinetic plus magnetic energy, while the implicit sourcing by numerical dissipation would encompass the integral effect of LES-like viscosity and resistivity terms entering the MHD equations (see, e.g., grete16; grete17; vlaykov16, for an example of such explicit subgrid MHD models). Note, however, that the validity of such models must be assessed with care as some of the basic assumptions of the LES framework—the scale locality and unidirectionality of the energy cascade—can be violated in the presence of magnetic fields (e.g., alexakis07; yousef07; stepanov15; grete17-mhdcascade; schekochihin22; friedrich24; beattie25). First, magnetic fields can enable nonlocal energy transport, breaking the assumption that local source terms for subgrid magnetized turbulence can depend solely on the local flow properties. Second, they enable an inverse cascade, in which energy flows from small, unresolved to large, resolved scales. In the context of the SLES model, the inverse cascade acts analogously to locally negative dissipation: the entropy of the resolved flow decreases as kinetic energy is transferred from unresolved to larger scales. Such negative dissipation is present in most second- and higher-order methods, but it is purely numerical in origin and must be treated with caution (see the discussion of entropy stability below). An MHD extension of the SLES model would therefore require an explicit treatment of the energy transfer behavior of the magnetic fields across the resolution scale, which can be particularly challenging for low-order methods, since such a model would need a large spatial kernel (>10>10 cells) to fully capture the dissipative scales where this transfer occurs numerically (see, e.g., grete17).
@@
The results of Section 4 show that the distribution of $e_{\rm turb}$ in the cold, supersonic ISM is only weakly sensitive to the choice of model. This is because in this regime the production of small-scale turbulence is dominated by the heating of preexisting turbulence via the $PdV$ work exerted by gas undergoing compression in spiral arms, collision of SN-driven shells, and disk instabilities (see semenov16; semenov17). Thus, the choice of the model for the cascade term, $\Sigma$, in Equation (1) might not be as crucial for predicting the turbulent properties of the cold ISM, as long as the adiabatic $PdV$ term and turbulence decay, $\epsilon$, are modeled appropriately. However, significant differences do arise in regions where the cascade of energy from resolved scales dominates the sourcing of $e_{\rm turb}$.
@@
One example is the disk outskirts and very center, where the SLES model predicts higher turbulent velocities than the explicit LES model (Section 4). One reason for this discrepancy is the difference in the treatment of shearing flows. Our explicit model adopts a shear-improved LES closure where the bulk flow is filtered out and only the fluctuating part of the velocity field is interpreted as a part of the turbulent cascade (leveque07, see also Section 2.2). To avoid spurious production of $e_{\rm turb}$ in stable shearing flows, such as differentially rotating disks, one needs to adopt a similar treatment in the SLES model. For example, instead of putting all dissipated energy into $e_{\rm turb}$, one can make it a function of the difference between the local filtered and total flows, such that in purely shearing flows, kinetic energy is dissipated directly into heat. We leave the investigation of such models to future work.